OpenStudy (goformit100):

The radii of the ends of a frustum of a cone 45 cm high are 28 cm and 7 cm (see Fig. 13.21). Find its volume, the curved s u r f a c e a r e a a n d t h e t o t a l s u f a c e a r e a

OpenStudy (goformit100):

@marimar @magenxoxo

OpenStudy (goformit100):

@timo86m

OpenStudy (goformit100):

@Jenn777 HELP

OpenStudy (goformit100):

@bngaure @nincompoop

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):



OpenStudy (anonymous):

now, R/H=r/(H-h) =>Hr=HR-hR =>hR=HR-Hr=H(R-r) =>H=Rh/(R-r)=28(45)/(28-7)=28(45)/21=4(45)/3=4(15)=60cm and, l=sqrt(r^2+h^2)=sqrt(7^2+45^2) L=sqrt(R^2+H^2)=sqrt(28^2+60^2) so, V=(1/3)[piR^2-pir^2]h =(pi*h/3)(R^2-r^2) Sc=pi(R-r)l St=Sc+pi(R^2+r^2) You evaluate yourself.